R is an ordered eld which properly contains Q as an ordered subf ield.

(a) Does R have elements 0 and 1 such that x+0 = x for all x 2 R and x  1 = x for

all x 2 R.

(b) Prove that 0 = 0 and 1 = 1 where 0 and 1 are the integers that we know very

well.


1
Expert's answer
2021-10-29T00:01:02-0400

Field is abelian group with respect to + and abelian group on R\{0} withrespext to * and "0 \ne1" . Therefore elements "1=1_R,0=0_R" exist.

Let "0_Q" zero in subfield Q. We have identity "0_Q+0_Q=0_Q"

Let "(-0_Q)" inverse in (R,+) element of "0_Q"

Tnen

"(-0_q)+0_Q+0_Q=(-0_Q)+0_Q;\\
0+0_Q=0;\\
0=0_Q"

By the same way "1_Q\cdot 1_Q=1_Q"

"1_Q\ne 0" because "0=0_Q"

Let  "1_Q^{-1}" be inverse of "1_Q" in R

Then

"1_Q^{-1}\cdot 1_Q\cdot 1_Q=1_Q^{-1}\cdot 1_Q=1=1_R"

"1_R^ \cdot 1_Q=1_R\\
1_Q=1_R"